OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (radar):

Subtract the two given equations getting -5y=5, or y=-1 that is all you need to do as there is only one choice that has y = -1, the third choice (12, -1)

OpenStudy (radar):

That was the lazy way, let me check and actually solve for x, substituting -1 for y in the first equation: x -2(-1)=14 x+2=14 x=14-2=12 Yup

OpenStudy (radar):

Any questions?

OpenStudy (anonymous):

that all thank u

OpenStudy (radar):

ur welcome and good luck in your studies.